OpenStudy (danmac0710):

Is it possible to write this equation explicitly in terms of y? xy/2 = x + y + (x^2+y^2)^(1/2)

OpenStudy (amistre64):

maybe the wolf can help out with that ...

OpenStudy (danmac0710):

Who or what is the wolf??

OpenStudy (danmac0710):

Wolfram Alpha!! Of course! Wow does it do that kind of thing? Impressive!

OpenStudy (asnaseer):

you should be able to do this quite simply by isolating the (x^2+y^2)^(1/2) on one-side of the equation and then squaring both sides.

OpenStudy (asnaseer):

\[\frac{xy}{2}=x+y+\sqrt{x^2+y^2}\]now multiply both sides by 2 to get:\[xy=2x+2x+2\sqrt{x^2+y^2}\]re-arrange to get:\[2\sqrt{x^2+y^2}=xy-2x-2y\]square both sides:\[4(x^2+y^2)=(xy-2x-2y)^2=x^2y^2-4x^2y+4x^2-4xy^2+8xy+4y^2\]the 4x^2 and 4y^2 appear on both sides of the equation and so can be cancelled out, leaving:\[0=x^2y^2-4x^2y-4xy^2+8xy\]you should be able to solve from here...